OpenStudy (anonymous):

\int\limits_0^1 x e^(-3x) dx

OpenStudy (raden):

integral by parts or by using tabular integration

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@amistre64

ganeshie8 (ganeshie8):

substitute \(-3x = z\)

OpenStudy (anonymous):

okay :)

OpenStudy (amistre64):

that still produces a by parts ....

ganeshie8 (ganeshie8):

why lol

OpenStudy (amistre64):

x = -z/3 dx = -3 dz

OpenStudy (amistre64):

well, got a little sloppy on my dx, but yeah

ganeshie8 (ganeshie8):

oops yeah my breain is dead... we need to work by parts here :o

OpenStudy (anonymous):

:) @amistre64 and :) @ganeshie8

ganeshie8 (ganeshie8):

familiar wid by parts ?

OpenStudy (anonymous):

okay i will try to do :)

ganeshie8 (ganeshie8):

try below : \(\large u = x^2/2 \) \(\large v = e^{-3x}\)

OpenStudy (anonymous):

thank you :)